ABSTRACT. A generating basis and the dual cone of n-convex functions satisfying certain constraints are derived. As applications, the existence and characterization of a best L_{p}-approximation (1 ≤ p < ∞) from such subcones to a function in L_{p} are established. The relationship between a best L_{1}-approximation and perfect splines is developed under certain conditions.